OpenStudy (anonymous):

Which is equal to 9^1/3 * 9^1/6

OpenStudy (the_fizicx99):

Do you want an answer or explanation?

OpenStudy (anonymous):

both

OpenStudy (the_fizicx99):

Mk, Do you know exponential rules? Basically \(\ \sf x^{\frac{1}{2}} \implies \sqrt{x} \) Therefore, \(\ \sf 9^{\frac{1}{3}} \implies \sqrt[3]{9} \) Applying it to both terms you get: \(\ \sf \sqrt[3]{9} \times \sqrt[6]{9}\) Can you simplify this?

OpenStudy (anonymous):

343.65?

OpenStudy (the_fizicx99):

In short yes, you can. The squareroot of 9 is 3^2. Therefore, \( \sf \sqrt[3]{3^2} \times \sqrt[6]{3^2} \) Now undoing the rule we applied, we result with \( \sf 3^{\frac{2}{3}} \times 3^{\frac{2}{6}} \) After simplfying you get 3^2/3 * 3^2/6 ==> 2/6 = 1/3 3^2/3 * 3^1/3; bases are the same, so add the exponents: 2/3 + 1/3 = 3/3 = 1 So 3^1 = 3. Your final answer is 3.

OpenStudy (the_fizicx99):

I have to go. Tag me if you need further help later on :)